OpenStudy (anonymous):

RSTU is a rectangle RX is congruent to YS Find RU and UX if UT=18 and the area of RSTU is 216u to the power of 2

OpenStudy (anonymous):



OpenStudy (anonymous):

RU=12 UX=13

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Area of rectangle RU * UT=216 RU * 18=216 RU=12 as we see triangle RXU is a right triangle so apply pythagores theorem RX^2+RU^2=UX^2 5^2+12^2=UX^2 169=UX^2 13^2=UX^2 UX=13

OpenStudy (anonymous):

ok thanks so much